A laser delivers pulses of 2mJ at wavelength 633 nm with a pulse duration of
200 ps. (hint E=h.c/l, with h=6.6 x10-34 J.s and c=3x108 m/s and 1 eV=1.6
x10-19 J)
i) How far do they travel in 1s? (5marks)
ii) What is the photon energy in eV?(10marks)
iii) How many photons are delivered per pulse?10 marks)
1
Expert's answer
2011-07-01T14:46:39-0400
i) S=c*t=3*108*1=3*108 meters.
ii) eV=1.6*10-19 J → J=1019/1.6 eV= 0.625*1019 eV=625*1016 eV → E=2*10-3*1016=2*1013 eV
iii) energy of 1 photon is E1 = hc/λ=6.6*10-34 *3*108/633*10-9 = 31.2*10-20 J
Thus
n = E/E1 = 2*10-3/31.2*10-20 = 6.4*31.2*1015 particles.